OpenStudy (anonymous):

Name the property illustrated in this statement. (8 + 17) + 3 = 8 + (17 + 3)

OpenStudy (anonymous):

which property do you think it is?

OpenStudy (anonymous):

Commutative property

jimthompson5910 (jim_thompson5910):

no, you're moving around parenthesis, so you're re-associating terms which means you're using the associative property of addition

OpenStudy (anonymous):

commutative property of addition commutative just changes around the number A+B = B+A associative property of addition associative property moves the parenthesis around (A+B) +C = A+ (B + C)